OpenStudy (yrelhan4):

Ok so I have a problem in logic somewhere. I have this question where a sphere is being pulled on a rough horizontal surface by a force. I have to find the angular acc when it starts pure rolling. Here I take the moment of inertia about the bottom point, that is in contact of the ground i.e. 7/5 M R^2 . Now another question where a sphere is rolling down a plane. The moment of inertia is considered about the centre point. Why so?

OpenStudy (yrelhan4):

@Shane_B @Mashy

OpenStudy (dls):

\[\LARGE \tau = I \times \alpha\]

OpenStudy (dls):

:P

OpenStudy (yrelhan4):

kya hai yaar. ye to mereko bhi pta hai. :/

OpenStudy (dls):

I=7/5MR^2 tau you can calculate,so what's the problem? :o

OpenStudy (yrelhan4):

i take 7/5MR^2 at one place and 2/5MR^2 at the other. thats the problem.

OpenStudy (dls):

okay let me think :| don't expect an answer though

OpenStudy (dls):

ah,still here ? D:

OpenStudy (yrelhan4):

haha yeah.

OpenStudy (dls):

I didn't encounter anything like this yet,where is the question from?

OpenStudy (yrelhan4):

the second one came in my eee last year. and the first one is in my book.

OpenStudy (dls):

ill tell you about second one,brb.just wait

OpenStudy (yrelhan4):

hey i have the solution. i got the right answer that time. i have a doubt about which point i should take the MOI.

OpenStudy (yrelhan4):

@Mashy

OpenStudy (anonymous):

how can you take M.I about the bottom point? .. its not rotating about that point!.. its rotating about the centre!

OpenStudy (dls):

until and unless its a pulley -_-

OpenStudy (dls):

thattoo stationary :P

OpenStudy (yrelhan4):

i dont know. instantaneous centre of rotation is what they call it.

OpenStudy (anonymous):

wat s that? :-/.. just take about the centre.. its easy!

OpenStudy (dls):

MOI about CM =0

OpenStudy (dls):

its IAOR

OpenStudy (dls):

instantaneous axis of rotation

OpenStudy (yrelhan4):

yes. what is that? @DLS

OpenStudy (anonymous):

ok its the point that is at rest.. when something is rolling without slip, the point in contact with ground is at rest.. so that is the instant centre or whatever!

OpenStudy (yrelhan4):

you dont take MOI about that point?

OpenStudy (dls):

An axis passing through IC and perpendicular to the plane on which the motion is being described is called IAOR

OpenStudy (yrelhan4):

so finally. do i take moi about iaor or the centre?

OpenStudy (dls):